OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(3\quad ,&-6) \end{array} \\\quad \\ slope = m= 4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values }\)

OpenStudy (jdoe0001):

\(\bf y-y_1=m(x-x_1)\) <-- point-slope form of an linear equation to get the slope-intercept form, just solve for "y" as far as the standard form -> http://www.mathsisfun.com/algebra/standard-form.html